# On the asymptotic behavior of the Repulsive Pressureless Euler-Poisson System

### v1

@@ 2.4 Instantaneous Failure of the Projection Formula
Before moving on, we will present an example in which (2.1) breaks instantaneously. Observe that the projection formula does not hold in the repulsive case for $t>t_{c}=2\sqrt{2}$ in Example 2.1. The goal here is to create a DSPS with nested situations like this, having $t_{c}$ arbitrarily small. Indeed, put a particle of mass $1/4$ at $y=1$, mass $1/8$ at $y=1/2$, mass $1/16$ at $y=1/4$, and so on. Mirroring this on the negatives, we have constructed $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R})$. The initial acceleration of each particle may be calculated by a geometric series. Then, give each particle just enough velocity to reach the origin, i.e., so that its parabolic trajectory has its vertex at y=0y=0.333See Definition 2.12 and section 3 for more on solutions like this. In all, we obtain the following initial conditions and solution:
@@
$\rho_{0}=\sum_{i=1}^{\infty}\frac{\delta_{3^{1-i}}+\delta_{-(3^{1-i})}}{2^{i+1}},\quad\quad v_{0}(3^{1-i})=-\operatorname{sgn}(y)\frac{3\sqrt{2}}{2}\big(\sqrt{6}\big)^{-i}$
@@ 2.5 Combining Particles in finite DSPS
The following section will be entirely in service of establishing Lemma 2.10, which will prove a useful tool in several future theorems. It shows that any contained group of particles can be replaced by a single particle at the center of mass of the group with the group’s total mass and total momentum, with no change to the relevant behavior of the solution.
@@
This is the initial parabolic trajectory extrapolated over all time. (i.e., the dotted lines in figure 2.1)
@@
For this and future proofs in this section, we will re-index for ease of notation: let i=1(q1)i=1-(q-1), $2-(q-1),\ldots,$ n(q1)n-(q-1), such that $p_{1}$ is the first particle of $P$. We proceed by induction on $l$, for which the $1$ particle base case is trivial. Now, assume this holds true for $l=k$, and we will show it for $l=k+1$. If there are no collisions within $P$ for $t<t_{P}$ the result is trivial, so we assume there is a collision within $P$ before $t_{P}$. Let $t_{c}$ be the time of first collision within $P$. We have the result trivially for $t\leq t_{c}$, so it remains to be shown for $t_{c}<t\leq t_{P}$. Let $p_{j}$ and $p_{j+1}$ be the particles involved in the first collision in $P$. (If there are multiple collisions at $t_{c}$, we can choose any adjacent pair which collides.) After collision, these two particles combine into a new particle $\bar{p}$ with mass $\bar{m}=m_{j}+m_{j+1}$. We can calculate its ghost trajectory for $t\geq t_{c}$, setting our initial condition to $t_{c}$ (using the semigroup property) and recalling that collisions observe conservation of momentum. Now considering the solution originating at $\tilde{\rho}_{0}=\rho_{t_{c}}$, we have $y_{\bar{p}}(0)=y_{j}(0)=y_{j+1}(0)$, $y^{\prime}_{\bar{p}}(0)=\frac{m_{j}y_{j}^{\prime}(t)+m_{j+1}y^{\prime}_{j+1}(t)}{m_{j}+m_{j+1}}$, and
@@
Using this lemma, we can begin to understand collisions using simplified initial conditions. This is the biggest step towards achieving the desired sort of particle combining which will be used in future proofs.
@@
It is clear from the definition that the trajectories of particles in $P$ bound $y_{p*}(t)$. Thus, by assumption particles in $P$ don’t collide with other particles before $t_{P}$, and since $p^{*}$ affects outside particles the same way as $P$, (as can be observed from (2.3)) we get that $p^{*}$ doesn’t collide with other particles for $t<t_{P}$, so the above calculation gives its trajectory for $t\leq t_{P}$ (we get $t=t_{P}$ by continuity of position). Now, recalling (2.8) and using Lemma 2.8, for $t\leq t_{P}$ we have
@@
Let us examine the solutions at $t=t_{Q}$. The solutions are identical for particles outside of $P$, since for all $t<t_{P}$ the mass in $P$ and $p^{*}$ are on the same side as every other mass in the solutions, between which the only initial differences occur in $P$. In $\rho$, $P$ has formed a single particle with mass equal to $m_{p^{*}}$ and trajectory $y_{P}(t)$, as the center of mass of the system is now just the position of the single particle. By Lemma 2.9, $y_{p^{*}}(t)=y_{P}(t)$ for all $t\leq t_{P}$, and both functions are differentiable on $(0,t_{P})$, so we have $y^{\prime}_{P}(t_{Q})=y^{\prime}_{p^{*}}(t_{Q})$, meaning that the newly formed particle has velocity at $t_{Q}$ equal to the velocity of $p^{*}$ at $t_{Q}$. Since $\rho$ and $\rho^{*}$ are identical mass distributions with (a.e.) identical velocity distributions at $t_{Q}$, existence and uniqueness of DSPS tells us that $\rho$ and $\rho^{*}$ must be identical for all times $t\geq t_{Q}$, as otherwise applying the semigroup property at $t_{Q}$ would yield two different solutions for the same initial condition.
@@
$\frac{\sum_{i=1}^{s+r}\tilde{m}_{i}\frac{d^{-}}{dt}\tilde{y}_{i}(t)}{\sum_{i=1}^{s+r}\tilde{m}_{i}}$
and, noting that by Lemma 2.8 we have
@@ 2.7 Strong Convergence of Discrete Approximations to GSPS
We know that $t\mapsto H(\rho_{t}^{n},v_{t}^{n})$, $t\in[0,T]$, is nonincreasing. Thus, we can establish that
@@ 3 Perfect Solutions
We now turn our attention to the primary result of the section: the existence and uniqueness of velocity distributions yielding perfect solutions given discrete initial mass distributions. Proof of this fact will require several lemmas, the first of which has much merit on its own to the understanding of glancing collisions. It shows that whether a first collision of two particles is glancing or not depends solely on the interplay between their masses, relative initial positions and velocities.
@@
We can now prove the main result of the section.
@@ 4.1 A Quadratic Bound on Compactly Supported Solutions
This bound is sharp; it traces out the trajectory of a two-particle perfect SPS, where the particles begin at $y_{L}(0)$ and $y_{R}(0)$, and the mass is distributed among them in order to make the center of mass 0.101010Compare this bound to the parabolic bound on evolution of any solution to the attractive PEP equations presented in Theorem 3.11 of [1]. One would expect that the concavity of the bounding parabolas is inverted; interestingly, this bound no longer depends on $\|v_{0}\|L^{\infty}(\rho_{0})$, as solutions with enough outward velocity will simply not converge to equilibrium.
@@ 6 Future Steps
The reverse direction is clear, but the forward direction we cannot prove because we do not know enough about the properties of the approximating solutions for any given GSPS. For example, the approximating DSPS may not converge to equilibrium at all, meaning that the above bound does not apply. However, the same intuition which is presented above should hold.

### v2

@@ 1.1 Overview
We note here that related works on biological aggregation deal with short range repulsive and long-range attractive interactions, variational solutions for, spatial confinement and the asymptotic behavior of the resulting dynamics. We mention only a few here: [1], [5], [7], [8].
@@ 2.4 Instantaneous Failure of the Projection Formula
Before moving on, we will present an example in which (2.1) breaks instantaneously. Observe that the projection formula does not hold in the repulsive case for $t>t_{c}=2\sqrt{2}$ in Example 2.1. The goal here is to create a DSPS with nested situations like this, having $t_{c}$ arbitrarily small. Indeed, put a particle of mass $1/4$ at $y=1$, mass $1/8$ at $y=1/3$, mass $1/16$ at $y=1/9$, and so on. Mirroring this on the negatives, we have constructed $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R})$. The initial acceleration of each particle may be calculated by a geometric series. Then, give each particle just enough velocity to reach the origin, i.e., so that its parabolic trajectory has its vertex at y=0y=0.333See Definition 2.12 and section 3 for more on solutions like this. In all, we obtain the following initial conditions and solution:
@@
$\rho_{0}=\sum_{i=1}^{\infty}\frac{\delta_{3^{1-i}}+\delta_{-(3^{1-i})}}{2^{i+1}},\quad\quad v_{0}(\pm 3^{1-i})=\mp\frac{3\sqrt{2}}{2}\big(\sqrt{6}\big)^{-i}$
@@ 2.5 Combining Particles in finite DSPS
The following section will be primarily in service of establishing Lemma 2.10, which will prove a useful tool in several future theorems. It shows that any contained group of particles can be replaced by a single particle at the center of mass of the group with the group’s total mass and total momentum, with no change to the relevant behavior of the solution.
@@
This is the initial parabolic trajectory extrapolated over all time (i.e., the dotted lines in figure 2.1).
@@
We will prove this by an examination of the first collision in $\rho$, allowing us to induct on the number of particles through an application of the semigroup property.
For this and future proofs in this section, we will re-index for ease of notation: let i=1(q1)i=1-(q-1), $2-(q-1),\ldots,$ n(q1)n-(q-1), such that $p_{1}$ is the first particle of $P$. We proceed by induction on $l$, for which the $1$ particle base case is trivial. Now, assume this holds true for $l=k$, and we will show it for $l=k+1$. If there are no collisions within $P$ for $t<t_{P}$ the result is trivial, so we assume there is a collision within $P$ before $t_{P}$. Let $t_{c}$ be the time of first collision within $P$. We have the result trivially for $t\leq t_{c}$, so it remains to be shown for $t_{c}<t\leq t_{P}$. Let $p_{j}$ and $p_{j+1}$ be a pair of particles involved in the first collision in $P$ (If there are multiple collisions at $t_{c}$, we can choose any adjacent pair which collides). [^4] One may complain that singling out two particles from the first collision in $P$ may not be accurate as, once the semigroup property is applied, the only composite particle which exists with its own trajectory is that of all particles involved in the first collision. We can address this with strong induction, reducing the system according to the first collision size. However, this argument has the exact same steps as the given, with far denser notation. For this reason, the two particle assumption is allowed. After collision, these two particles combine into a new particle $\bar{p}$ with mass $\bar{m}=m_{j}+m_{j+1}$. We can calculate its ghost trajectory for $t\geq t_{c}$, setting our initial condition to $t_{c}$ (using the semigroup property) and recalling that collisions observe conservation of momentum. Now considering the solution originating at $\tilde{\rho}_{0}=\rho_{t_{c}}$, we have $y_{\bar{p}}(0)=y_{j}(0)=y_{j+1}(0)$, $y^{\prime}_{\bar{p}}(0)=\frac{m_{j}y_{j}^{\prime}(t)+m_{j+1}y^{\prime}_{j+1}(t)}{m_{j}+m_{j+1}}$, and
@@
We can now begin to understand collisions using simplified initial conditions. Using the previous lemma, we now prove that because the group’s center of mass is equal to the center of mass of the ghost trajectories, it is also equal (at least until $t_{P}$) to the trajectory of a single “center of mass” particle which replaces the entire group from the start.
@@
It is clear from the definition that the trajectories of particles in $P$ bound $y_{p*}(t)$. Thus, by the assumption that particles in $P$ don’t collide with other particles before $t_{P}$, and since $p^{*}$ affects outside particles the same way as $P$ (as can be observed from (2.3)), we get that $p^{*}$ doesn’t collide with other particles for $t<t_{P}$, so the above trajectory holds for $t\leq t_{P}$ (we get $t=t_{P}$ by continuity of position). Now, recalling (2.8) and using Lemma 2.8, for $t\leq t_{P}$ we have
@@
We are now ready to prove the primary result of this section. The following lemma demonstrates that the replacement used in the previous lemma leaves the rest of the solution unchanged, and that when the particles in $P$ all combine at $t_{P}$, the particle they form has the same trajectory as that of $p^{*}$ at $t_{P}$, meaning that the combined and uncombined solutions proceed identically following $t_{P}$ (and therefore they are equivalent in terms of equilibrium).
@@
Let us examine the solutions at $t=t_{Q}$. The solutions are identical for particles outside of $P$, since for all $t<t_{P}$ the mass in $P$ and $p^{*}$ are equal and are on the same side of every other mass in the solutions. In $\rho$, $P$ has formed a single particle with mass equal to $m_{p^{*}}$ and trajectory $y_{P}(t)$, as the center of mass of the system is now just the position of the single particle. By Lemma 2.9, $y_{p^{*}}(t)=y_{P}(t)$ for all $t\leq t_{P}$, and both functions are differentiable on $(0,t_{P})$, so we have $y^{\prime}_{P}(t_{Q})=y^{\prime}_{p^{*}}(t_{Q})$, meaning that the newly formed particle has velocity at $t_{Q}$ equal to the velocity of $p^{*}$ at $t_{Q}$. Since $\rho$ and $\rho^{*}$ are identical mass distributions with (a.e.) identical velocity distributions at $t_{Q}$, existence and uniqueness of DSPS tells us that $\rho$ and $\rho^{*}$ must be identical for all times $t\geq t_{Q}$, as otherwise applying the semigroup property at $t_{Q}$ would yield two different solutions for the same initial condition.
@@
$\frac{\sum_{i=1}^{s+r}\tilde{m}_{i}\frac{d^{-}}{dt}\tilde{y}_{i}(t)}{\sum_{i=1}^{s+r}\tilde{m}_{i}}.$
Now, noting that by Lemma 2.8 we have
@@ 2.7 Strong Convergence of Discrete Approximations to GSPS
Given our wealth of tools for studying discrete measures, it is natural that we should want some mechanism to extend discrete results to more general measures. The following theorem gives the conditions under which a series of discrete approximations converge appropriately to a more general measure. The proof makes use of our results on the Hamiltonian, as well as crucial results from Proposition 3.3 in [14], in order to extract appropriate bounds, approximations, and subsequences to obtain the result.
@@
By Theorem 2.17, $t\mapsto H(\rho_{t}^{n},v_{t}^{n})$ is nonincreasing for $t\in[0,T]$. Thus, we can establish that
@@ 3 Perfect Solutions
We now turn our attention to the primary result of the section: the existence and uniqueness of velocity distributions yielding perfect solutions given discrete initial mass distributions. Proof of this fact will require several lemmas, the first of which has much merit on its own to the understanding of glancing collisions. It shows that whether a first collision of two particles is glancing depends solely on the interplay between their masses and their relative initial positions and velocities, or equivalently the time of their collision. The argument is mainly algebraic.
@@
The following lemma handles a specific situation relevant to Theorem 3.4, and is the final preliminary to its proof.
@@
We can now prove the main result of the section, by showing that in a perfect solution the pair of particles which minimize $t_{c}$ given in Lemma 3.2 must be the first pair to collide. This allows us to induct on the number of particles by applying results from Section 2.5 to this first collision.
@@ 4.1 A Quadratic Bound on Compactly Supported Solutions
This bound is sharp; it traces out the trajectory of a two-particle perfect SPS, where the particles begin at $y_{L}(0)$ and $y_{R}(0)$, and the mass is distributed among them in order to make the center of mass 0.111111Compare this bound to the parabolic bound on evolution of any solution to the attractive PEP equations presented in Theorem 3.11 of [2]. One would expect that the concavity of the bounding parabolas is inverted; interestingly, this bound no longer depends on $\|v_{0}\|_{L^{\infty}(\rho_{0})}$, as solutions with enough outward velocity will simply not converge to equilibrium.
@@ 6 Future Steps
The reverse implication is clear. As for the direct one, we would need to prove existence of approximating DSPS which converge to equilibrium so that the envelope (4.1) applies uniformly; then we can pass to the limit and make the desired inference on the GSPS. Even if one can prove uniform bounds on the Wasserstein distance $W_{2}(\rho_{t}^{n},\rho_{t})$, the fact that $\rho_{t}=\delta_{0}$ for sufficiently large $t$ would not suffice to infer that for sufficiently large $n$ the quantity $W_{2}(\rho_{t}^{n},\rho_{t})$ converges to zero as $t\rightarrow\infty$. Thus, it is not clear how to apply Theorem 4.1 in order to prove the above conjecture.
@@
Given the nonincreasing behavior of the Hamiltonian, we do not expect rapid oscillations around $\delta_{0}$; in fact, if the GSPS collapses to $\delta_{0}$ in finite time, then the velocity trivially vanishes from the collapse time forward.